OpenStudy (anonymous):

write the equation given these characteristics in a parabola : focus (-3,-2) and directrix y=1

OpenStudy (anonymous):

formula should b (x-h)^2 = 4p (y-k) I know the vertex will be (-3, _____) i dont get how to find the y in the vertex and the p

OpenStudy (anonymous):

well u r missing some equations for the focus and the directrix

OpenStudy (anonymous):

for example for the focus u have (h, k+p) =(-3,-2) and for the directrix u have y=1=k-p

OpenStudy (anonymous):

but what is k ? how did you find it

OpenStudy (anonymous):

u have 2 equation with 2 variales that is k+p = -2 1= k-p

OpenStudy (anonymous):

i get that but how did you get the value for each it is a bit confusing

OpenStudy (anonymous):

vortex is always halfway between diretrix and focus so vertex must be (-3,-0.5) so h=-3 k=-0.5 p=(distance from focus to vortex ) p=1.5

OpenStudy (anonymous):

did you get it ?

OpenStudy (anonymous):

thankssss i did